OpenStudy (anonymous):

Ashley is reading the blueprint above to build a game room. The scale is 1 cm = 10 feet. If the length of the south wall is 40 feet, how many centimeters is y? 4 cm

OpenStudy (pfenn1):

This is a ratio problem. If 1 cm on plan = 10 feet in life, then ?cm=40 ft.\[\frac{1}{10}=\frac{x}{40}\]

OpenStudy (anonymous):

cant get any simpler huh thx

OpenStudy (anonymous):

I prefer dimensional analysis 40 ft * 1 cm / 10 ft ( conversion factor) the ft cancels out :)

OpenStudy (anonymous):

left only with cm

OpenStudy (pfenn1):

They are really the same thing. \[\frac{1 cm}{10 ft}=\frac{x cm}{40 feet}\]Rearrange\[x=\frac{40\cancel{ft}}{10\cancel{ft}}\times 1cm\] Same exact thing.